# Fully Energy-Efficient Randomized Backoff: Slow Feedback Loops Yield Fast Contention Resolution

### v2

@@ Abstract.
A shared communication channel can also suffer random or adversarial noise, which causes any listener to hear noise, even when no packets are actually sent. Such noise arises due to hardware/software failures or malicious interference (all modeled as “jamming”), which can have a ruinous effect on the throughput and energy efficiency. How does noise affect our goal of long feedback loops/energy efficiency?
@@
Tying these questions together, we ask: what does a contention-resolution algorithm have to sacrifice to reduce channel accesses? Must we give up on constant throughput? What about robustness to noise? Here, we show that we need not concede anything by presenting an algorithm with the following guarantees. Suppose there are $N$ packets arriving over time and $\mathcal{J}$ jammed slots, where the input is determined by an adaptive adversary. With high probability in $N+\mathcal{J}$, our algorithm guarantees $\Theta(1)$ throughput and $\operatorname{polylog}(N+\mathcal{J})$ channel accesses (sends or listens) per packet. We also have analogous guarantees when the input stream is infinite.
1. Introduction
@@
Since the 1970s, randomized backoff protocols (Goldberg-notes-2000, ; HastadLeRo96, ; MetcalfeBo76, ; AbramsonKu73, ; 802.11-standard, ), such as binary exponential backoff (MetcalfeBo76, ), have been used for managing contention on a shared communication channel. Originally used in the ALOHA system (ALOHAnet, ; binder:aloha, ) and Ethernet (MetcalfeBo76, ; stuck1984introduction, ), randomized backoff plays an important role in a wide range of applications, including WiFi (802.11-standard, ), wireless sensor networks (khanafer2013survey, ), cellular systems  (cellular-standard, ), transactional memory (scherer2005advanced, ; HerlihyMo93, ), and congestion control (wierman2003unified, ). The salient feature of the communication channel is that it supports only one message transmission at a time: if more than one message is sent simultaneously, there is a collision resulting in indecipherable noise (Goldberg-notes-2000, ; kurose:computer, ; xiao:performance, ; DBLP:journals/jacm/GoldbergMPS00, ; MetcalfeBo76, ).
@@
This contention-resolution problem is formalized as follows (hastad:analysis, ; goldberg2004bound, ; kelly1987number, ; aldous:ultimate, ; GoodmanGrMaMa88, ; mosely1985class, ; RaghavanU98, ; chlebus2009maximum, ; BenderGiKu22, ; zhou2019singletons, ; BenderFiGi19, ; FinemanNW16:multiple, ; fineman:contention, ; DeMarcoSt17, ; BenderFiGi16, ; FernandezAntaMoMu13, ; BenderFaHe05, ; deMarco:fast, ; DEMARCO20171, ; ChangJiPe19, ; BenderKoKu20, ; BenderKoPe16, ). There are $N$ packets arriving over time, and each packet needs to be successfully transmitted on the channel. [^1] For ease of exposition, we slightly abuse terminology and have the packets themselves taking action (e.g., sending themselves on the channel, listening on the channel), rather than introducing “agents”/“devices”/“senders”, where each one arrives on the scene with a packet to transmit. Time is divided into synchronized slots, each of which is sized to fit a single packet. To be successfully sent, the packet requires exclusive access to the channel (ChlamtacKu85, ; Goldberg-notes-2000, ; xiao:performance, ; DBLP:journals/jacm/GoldbergMPS00, ; hastad:analysis, ; MetcalfeBo76, ; AbramsonKu73, ); that is, the packet must be the only one transmitted during that slot. Otherwise, if two or more packets are transmitted in the same slot, the result is a collision, where none of the transmitted packets succeed. A packet departs the system once it is successfully sent. The objective is to send all packets while optimizing the throughput of the channel, which is roughly the fraction of successful slots.
@@
The most famous contention-resolution protocol is binary exponential backoff (MetcalfeBo76, ) (see (kelly:decentralized, ; kelly1987number, ; rosenkrantz:some, ; aldous:ultimate, ; GoodmanGrMaMa88, ; GoodmanGrMaMa88, ; Al-Ammal2000, ; al2001improved, ; song:stability, )). Informally, under binary exponential backoff, a packet that has been in the system for $t$ time slots is sent with probability $\Theta(1/t)$.
@@
An elegant, but ultimately problematic, feature of classical binary exponential backoff is that it is oblivious—packets remaining in the system do not use channel feedback to adjust their behavior: a packet with age $t$ sends with probability $\Theta(1/t)$ until it succeeds, regardless of channel history. The unfortunate result is that with adversarial packet arrivals, [^2] There is also much work analyzing the stochastic arrival rates under which binary exponential backoff is stable or unstable (HastadLeRo87, ; aldous:ultimate, ; GoodmanGrMaMa88, ); see also (MetcalfeBo76, ; Goldberg-notes-2000, ; hastad:analysis, ; goldberg2004bound, ; shoch1980measured, ; abramson1977throughput, ; kelly1987number, ; aldous:ultimate, ; GoodmanGrMaMa88, ; mosely1985class, ; RaghavanU98, ; RaghavanU95, ; GOLDBERG1999232, ). binary exponential backoff supports only a subconstant throughput—specifically, $O(1/\log{N})$; in fact, even for the batch case where all $N$ packets arrive at the same time, the throughput of binary exponential backoff is only $O(1/\log N)$ (BenderFaHe05, ).
In contrast, contention-resolution protocols can exploit frequent channel feedback to achieve $\Theta(1)$ throughput under adversarial arrivals (BenderKoKu20, ; BenderFiGi19, ; BenderFiGi16, ; BenderKoPe18, ; BenderKoPe16, ; ChangJiPe19, ; awerbuch:jamming, ). These protocols are not oblivious—packets listen on the channel and adjust their sending probabilities up or down based on this feedback.
The primary model for feedback on the channel is the ternary-feedback model (richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; BenderFiGi19, ; DBLP:conf/spaa/AgrawalBFGY20, ; BenderFiGi16, ; BenderKoPe18, ; gilbert:near, ; DBLP:journals/dc/KingPSY18, ; chlebus2009maximum, ; chlebus2007stability, ; ChlebusKoRo06, ; fineman:contention2, ; anantharamu2019packet, ; greenberg1985lower, ). In this model, a packet can listen on the channel in each slot and learn whether that slot is (0) empty, if no packets send, (1) successful, if exactly one packet sends, or (2+) noisy, if two or more packets send. Based on this ternary feedback, the packet can decide when to attempt to send next. Once a packet succeeds, it immediately departs the system.
Most constant-throughput algorithms with ternary-feedback (e.g., (BenderKoKu20, ; BenderFiGi19, ; BenderFiGi16, ; BenderKoPe16, ; ChangJiPe19, ; DeMarcoSt17, ; awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming4, ; richa:jamming3, ; richa:efficient-j, ; ogierman:competitive, )) listen on the channel in every slot (or every constant number of slots). That is, these algorithms have a short feedback loop: a packet’s sending probability in slot $t+1$ is fully determined by its internal state at slot $t$ and the channel feedback at slot $t$. For example, the algorithm by Chang, Jin, and Pettie (ChangJiPe19, ), listens in every slot $t$ and multiplicatively updates the sending probability in slot $t+1$ based on whether it heard silence, a successful transmission, or a collision in slot $t$.
@@
Robustness and noise.
Finally, in addition to the factors discussed above, much of the recent work on contention resolution (anantharamu2019packet, ; anantharamu2011medium, ; Chlebus:2016:SWM:2882263.2882514, ; DBLP:conf/spaa/AgrawalBFGY20, ; ChenJiZh21, ; jiang:robust, ; awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming4, ; richa:jamming3, ; richa:efficient-j, ; ogierman:competitive, ; BenderFiGi19, ; BenderFiGi16, ; ChangJiPe19, ) has sought to address an additional factor, which is that the real world is often noisy. Sometimes interference prevents transmissions in a slot and listeners hear noise even if nobody actually sends. Noisy channels arise due to hardware/software failures, co-located devices, or malicious jamming (chiwewe2015using, ; liu2017interference, ; pelechrinis:denial, ; mpitziopoulos:survey, ; xu:feasibility, ). Regardless of the source of the noise in a slot, we may think of these slots as being “jammed” by an adversary. Jamming has evolved from a mostly-theoretical risk into a credible threat to systems over the past decade, with several publicized examples (nicholl, ; FCC:humphreys, ; marriot, ; priest, ).
After a long line of work, there are now many contention-resolution protocols that achieve constant throughput in the presence of noise (e.g. (awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming4, ; richa:jamming3, ; richa:efficient-j, ; ogierman:competitive, ; BenderFiGi19, ; BenderFiGi16, ; ChangJiPe19, )); however, again, these protocols listen to the channel in every slot and update their sending probabilities accordingly. Our goal is to eliminate short feedback loops not just in the classical model, but also with jamming.
@@
Up until now, we have discussed whether one can minimize listening and thus avoid providing immediate feedback for a contention-resolution algorithm. Another way of viewing this problem is through the lens of energy efficiency. Each channel access—whether for sending or listening—consumes energy. Most work on contention resolution is sending efficient, but is not listening efficient (e.g.,  (awerbuch:jamming, ; richa:jamming2, ; richa:competitive-j, ; richa:jamming3, ; richa:efficient-j, ; richa:jamming4, ; BenderFiGi19, ; BenderKoKu20, ; DeMarcoSt17, ; jiang:robust, ; de2022contention, )). That is, most protocols optimize how frequently a packet sends on the channel, but allow a packet to listen on the channel in every slot “for free”.
In fact, both sending and listening are expensive operations (e.g., (feeney2001investigating, ; polastre:telos, ; yang2017narrowband, ; lauridsen2018empirical, )), and minimizing energy usage by having devices sleep as much as possible has been a long-standing and popular strategy to maximize the network lifetime (e.g., (merlin2010duty, ; ye2006ultra, ; li2005energy, ; deng2005scheduling, )).333 Why bother optimizing listening in contention resolution protocols if packets must listen to receive messages (for other network applications)? Informally, there are two types of listening: listening to receive messages, and listening to execute a contention resolution protocol (which is what enables sending messages). In order to receive messages, packets do not need to listen in every time slot; there exist methods for optimizing this first type of listening. Although this is beyond the scope of this paper, as one example, in many wireless settings, a central base station can monitor the channel constantly and facilitate message exchange (feng2012survey, ). We call a protocol fully-energy efficient if it is both sending efficient and listening efficient. By definition, such a protocol cannot have short feedback loops, since it can access the channel only rarely.
@@
Currently, the only known path to full energy efficiency is via explicit synchronization. This means that the model is extended so that packets can send synchronization messages to each other whenever they broadcast (BenderKoPe18, ; DeMarcoSt17, ; marco:time, ). In (BenderKoPe18, ), these synchronization messages have size $\Theta(\log N)$ bits each, which means that an arbitrary polynomial amount of communication can be expressed in a slot. In (DeMarcoSt17, ; marco:time, ), the synchronization messages are smaller ($O(1)$ bits), but some packets are permitted to stick around as “Good Samaritans” in order to serve as long-term coordinators (that send many messages via many broadcasts).
Using $\Theta(\log N)$-bit synchronization messages, Bender, Kopelowitz, Pettie, and Young (BenderKoPe18, ) give an algorithm with $\Theta(1)$ throughput and expected $O(\log(\log^{*}N))$ channel accesses per packet. Using Good-Samaritan packets, De Marco and Stachowiak (DeMarcoSt17, ) and De Marco, Kowalski, and Stachowiak (marco:time, ) provide a constant-throughput algorithm that is sending efficient ($O(\log N)$ transmissions per packet) and conjecture that their techniques can be extended to achieve fully energy efficiency ($O(\operatorname{polylog}N)$ channel accesses per packet), with high probability and even without collision detection.
@@
We show that it is indeed possible to achieve constant throughput with full energy efficiency (i.e., both listening efficiency and sending efficiency) while supporting robustness to adversarial noise. Moreover, these results hold in the standard ternary-feedback model, without requiring the addition of any sort of explicit synchronization.
@@
Our algorithm belongs to a natural family of multiplicative-weight-update algorithms (e.g., (ChangJiPe19, ; awerbuch:jamming, ; richa:jamming2, ; richa:jamming3, ; ogierman:competitive, ; richa:jamming4, )). When a packet hears silence, it multiplicatively increases both its listening and sending probabilities. Conversely, when a packet hears noise, it multiplicatively decreases these probabilities. There are no control messages, no Good Samaritan packets, and no leaders elected.
@@
What makes our algorithm/analysis interesting is that we are able to support a multiplicative-weight-update framework while having each packet ‘cover its eyes’ almost all of the time. This is in stark contrast to prior constant-throughput algorithms, which adjust sending probabilities in every slot. Because each packet listens to so few slots, the different packets that are in the system at the same time may end up with very different perspectives on the world from each other. In order to analyze the ‘herd behavior’ of the packets in this potentially chaotic setting, substantially new techniques end up being required. These techniques are also what allow us to handle the additional chaos that adversarial jamming (i.e., noise) further adds to the system.
1.1. Model
@@
A finite or infinite stream of indistinguishable packets arrives over time; the number of arrivals is unknown to the algorithm. Each packet needs to be sent on the multiple-access channel. Time is divided into synchronized slots, each of which is sufficiently large to send a single packet. An adversary (specified below) controls precisely how many packets are injected into the system in each slot. When a packet successfully transmits, it leaves the system. There is no universal numbering scheme for the slots; that is, there is no global clock from which a newly injected packet could infer the lifetime of the system or slot parity. Additionally, the packets do not receive any additional information about how many packets have arrived or will arrive. Instead, packets only receive information through the ternary feedback model.
@@
We now describe the ternary feedback model (richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; BenderFiGi19, ; DBLP:conf/spaa/AgrawalBFGY20, ; BenderFiGi16, ; BenderKoPe18, ; gilbert:near, ; DBLP:journals/dc/KingPSY18, ; chlebus2009maximum, ; chlebus2007stability, ; ChlebusKoRo06, ; fineman:contention2, ; anantharamu2019packet, ; greenberg1985lower, ; willard:loglog, ). Initially, we will define the model without jamming. In each time slot, each packet in the system can take one of three actions: (i) sleep, (ii) send, or (iii) listen to the channel. Packets that take actions (ii) or (iii) are said to access the channel. If no packets choose to send during a slot, then that (non-jammed) slot is empty/silent; if exactly one packet sends, then that (non-jammed) slot is full and successful; if two or more packets send, then that slot is full and noisy. A packet that listens during a slot (action iii) learns whether the slot was (0) empty, (1) successful, or (2+) noisy. A packet that sleeps during a slot (action i) learns nothing about the state of the slot. A packet that sends (action ii), either transmits successfully and leaves the system, or collides and remains in the system. [^4] For ease of presentation, in our algorithms, we say that a packet can listen and send simultaneously, but any packet that is sending actually does not need to listen to determine the state of the channel. If the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.
@@
The basic metric: (overall) throughput. The main objective of contention resolution is to optimize throughput, defined next. A slot is active if at least one packet is in the system during that slot; inactive slots can be completely ignored in our analysis. Without loss of generality, assume throughout that the first slot is active. Without jamming, the throughput at time $t$ is defined as $T_{t}/S_{t}$, where $T_{t}$ is the number of successful transmissions during slots $1,2,\ldots,t$, and $S_{t}$ is the number of active slots in the same time interval. [^5] By assumption that the first slot is active, we have $S_{t}\geq 1$ and hence throughput is always well defined. On a finite input, the (overall) throughput is defined with respect to the final active slot $t$; at this point, the overall throughput is $N/S$, where $N=T_{t}$ is the total number of packets and $S=S_{t}$ is the total number of active slots. Overall throughput is not well-defined on an infinite execution.
A deficiency of the throughput metric (when defined naïvely) is that even if an algorithm guarantees $\Theta(1)$ overall throughput, it is not possible to achieve $\Theta(1)$ throughput uniformly across time. For example, if there is a burst of $N$ packets at time 00, and $N$ is unknown to the algorithm, then the throughput will be 00 for a superconstant number of slots (e.g., see (willard:loglog, ; ChangKoPe19, )), and this is provably unavoidable regardless of the backoff strategy being used.
Thus (overall) throughput is only meaningful at the end of the execution, or at points in time where there are no packets in the system.
A stronger metric: implicit throughput (BenderKoKu20, ). In order to support a meaningful notion of throughput, even at intermediate points in time, Bender, Kopelowitz, Kuszmaul, and Pettie (BenderKoKu20, ) propose a refined definition that they term “implicit throughput” (BenderKoKu20, ). The implicit throughput at time $t$ is defined as $N_{t}/S_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$ and $S_{t}$ is the total number of active slots so far. [^6] Notice that while $N_{t}$ depends on the adversary, the number of active slots depends on the contention resolution algorithm, whose goal it is to make slots inactive by completing packets.
@@
Throughout the rest of the paper, we will focus exclusively on implicit throughput. We should emphasize, however, that this only makes our results stronger—the results also imply the standard constant-throughput guarantees that one would normally strive for.
Extending to adversarial jamming. We next extend the definitions of throughput and implicit throughput for the case of jamming following (BenderFiGi16, ; BenderFiGi19, ). An algorithm wastes a slot if that slot has silence or a collision, and throughput measures the fraction of slots that the algorithm could have used but instead wasted. Let $\mathcal{J}_{t}$ denote the number of jammed slots through slot $t$. Then, the throughput of an execution ending at time $t$ is defined to be $(T_{t}+\mathcal{J}_{t})/S_{t}$, and the implicit throughput at slot $t$ is defined to be $(N_{t}+\mathcal{J}_{t})/S_{t}$.
@@
Observation 2 ( (BenderKoKu20, )).
Let $\delta$ be any lower bound on the implicit throughput of an algorithm; that is, suppose the algorithm achieves implicit throughput of at least $\delta$ at all times. Let $N_{t}$ and $S_{t}$ denote the total number of packet arrivals and active slots, respectively, at or before time $t$. Then $S_{t}\leq N_{t}/\delta$. Consequently

### v5

@@ Abstract
A shared channel can also suffer random or adversarial noise, which causes any listener to hear noise, even when no packets are actually sent. Such noise arises due to hardware/software failures or malicious interference (all modeled as “jamming”), which can have a ruinous effect on the throughput and energy efficiency. How does noise affect our goal of long feedback loops/energy efficiency?
@@
Tying these questions together, we ask: what does a contention-resolution algorithm have to sacrifice to reduce channel accesses? Must we give up on constant throughput? What about robustness to noise? Here, we show that we need not concede anything by presenting an algorithm with the following guarantees. Suppose there are $N$ packets arriving over time and $\mathcal{J}$ jammed slots, where the input is determined by an adaptive adversary. With high probability in $N+\mathcal{J}$, our algorithm guarantees $\Theta(1)$ throughput and $\operatorname{polylog}(N+\mathcal{J})$ channel accesses (sends or listens) per packet. We also have analogous guarantees when the input stream is infinite—we prove implicit throughput bounds of $\Omega(1)$ for all time slots $t$, and this translates to $\Theta(1)$ guaranteed throughput for any slot $t$ where the implicit throughput is sufficiently small in $\Theta(1)$. As a special case, these throughput results give rise to adversarial-queuing theory guarantees.
1 Introduction
@@
Since the 1970s, randomized backoff protocols such as binary exponential backoff [MetcalfeBo76], have been used for managing contention on a shared communication channel. Originally used in the ALOHA system [ALOHAnet] and Ethernet [MetcalfeBo76], randomized backoff plays an important role in a wide range of applications, including WiFi [802.11-standard], wireless sensor networks [khanafer2013survey], transactional memory [scherer2005advanced, HerlihyMo93], and congestion control [wierman2003unified]. The salient feature of the communication channel is that it supports only one message transmission at a time: if more than one message is sent simultaneously, there is a collision resulting in indecipherable noise [Goldberg-notes-2000, kurose:computer, xiao:performance, DBLP:journals/jacm/GoldbergMPS00, MetcalfeBo76].
@@
This contention-resolution problem is formalized as follows. There are $N$ packets arriving over time, and each packet needs to succeed (be the only packet sending) on the channel. [^1] For ease of exposition, we slightly abuse terminology and have the packets themselves taking action (e.g., sending themselves on the channel, listening on the channel), rather than introducing “agents”/“devices”/“senders”, where each one arrives on the scene with a packet to transmit. Time is divided into synchronized slots, each of which is sized to fit a single packet. To succeed, the packet requires exclusive access to the channel that is, the packet must be the only one transmitted during that slot. Otherwise, if two or more packets are transmitted in the same slot, the result is a collision, where none of the transmitted packets succeed. A packet departs the system once it succeeds. There is no a priori coordinator or central authority; packet transmissions are scheduled in a distributed manner. The objective is to have all packets succeed while optimizing the throughput of the channel, which is roughly the fraction of successful slots.
@@
A popular contention-resolution protocol is binary exponential backoff [MetcalfeBo76] (see [kelly:decentralized, kelly1987number, rosenkrantz:some, aldous:ultimate, GoodmanGrMaMa88, GoodmanGrMaMa88, Al-Ammal2000, al2001improved, song:stability]). Informally, under binary exponential backoff, a packet that has been in the system for $t$ slots is sent with probability $\Theta(1/t)$.
@@
An elegant, but ultimately problematic, feature of classical binary exponential backoff is that it is oblivious—packets remaining in the system do not use channel feedback to adjust their behavior: a packet with age $t$ sends with probability $\Theta(1/t)$ until it succeeds, regardless of channel history. The unfortunate result is that with adversarial packet arrivals, binary exponential backoff supports only a subconstant throughput—specifically, $O(1/\ln{N})$; in fact, even for the batch case where all $N$ packets arrive at the same time, the throughput of binary exponential backoff is only $O(1/\ln N)$ [BenderFaHe05].
In contrast, contention-resolution protocols can exploit frequent channel feedback to achieve $\Theta(1)$ throughput under adversarial arrivals [BenderKoKu20, BenderFiGi19, BenderKoPe16, ChangJiPe19, awerbuch:jamming]. These protocols are not oblivious—packets listen on the channel and adjust their sending probabilities up or down based on this feedback.
The primary model for channel feedback is the ternary-feedback model [BenderFiGi19, DBLP:conf/spaa/AgrawalBFGY20, gilbert:near, DBLP:journals/dc/KingPSY18, chlebus2009maximum, chlebus2007stability, ChlebusKoRo06, fineman:contention2, anantharamu2019packet, greenberg1985lower]. In this model, a packet can listen on the channel in each slot and learn whether that slot is (0) empty, if no packets send, (1) successful, if exactly one packet sends, or (2+) noisy, if two or more packets send. Based on this feedback, the packet can decide when to attempt to send next. Once a packet succeeds, it immediately departs the system.
Most constant-throughput algorithms with ternary-feedback (e.g., [BenderKoKu20, BenderFiGi19, BenderKoPe16, ChangJiPe19, DeMarcoSt17, awerbuch:jamming, richa:jamming2, richa:jamming4, richa:jamming3, ogierman:competitive]) listen on the channel in every slot (or every constant number of slots). That is, these algorithms have a short feedback loop: a packet’s sending probability in slot $t+1$ is fully determined by its internal state at slot $t$ and the channel feedback at slot $t$. For example, the algorithm by Chang, Jin, and Pettie [ChangJiPe19], listens in every slot $t$ and multiplicatively updates the sending probability in slot $t+1$ based on whether it heard silence, a successful transmission, or a collision in slot $t$.
@@
Robustness to noise.
Finally, in addition to the factors discussed above, much of the recent work on contention resolution (see [anantharamu2019packet, Chlebus:2016:SWM:2882263.2882514, DBLP:conf/spaa/AgrawalBFGY20, ChenJiZh21, jiang:robust, awerbuch:jamming, richa:jamming2, richa:jamming4, richa:jamming3, ogierman:competitive, BenderFiGi19, ChangJiPe19]) has sought to address an additional factor, which is that the real world is often noisy. Sometimes interference prevents transmissions in a slot and listeners hear noise even if nobody actually sends. Noisy channels arise due to hardware/software failures, co-located devices, or malicious jamming [chiwewe2015using, liu2017interference, pelechrinis:denial, mpitziopoulos:survey, xu:feasibility]. Regardless of the source of the noise in a slot, we may think of these slots as being “jammed” by an adversary. Jamming has evolved from a mostly-theoretical risk into a credible threat to systems over the past decade, with several publicized examples [nicholl, FCC:humphreys, marriot, priest].
After a long line of work, there are now many contention-resolution protocols that achieve constant throughput in the presence of noise (e.g. [awerbuch:jamming, richa:jamming2, richa:jamming4, richa:jamming3, ogierman:competitive, BenderFiGi19, ChangJiPe19]); however, again, these protocols listen to the channel in every slot and update their sending probabilities accordingly. Our goal is to eliminate short feedback loops not just in the classical model, but also with jamming.
@@
Up until now, we have discussed whether one can minimize listening and thus avoid providing immediate feedback for a contention-resolution algorithm. Another way of viewing this problem is through the lens of energy efficiency. Each channel access—whether for sending or listening—consumes energy. Most work on contention resolution is sending efficient, but is not listening efficient (e.g.,  [awerbuch:jamming, richa:jamming2, richa:jamming3, richa:jamming4, BenderFiGi19, BenderKoKu20, DeMarcoSt17, jiang:robust, de2022contention]). That is, most protocols optimize how frequently a packet sends, but allow a packet to listen in every slot “for free”.
In fact, both sending and listening are expensive operations (e.g., [polastre:telos, feeney2001investigating, yang2017narrowband, lauridsen2018empirical]), which should be minimized if devices are energy-constrained (e.g., battery-powered devices). Minimizing energy usage by having devices sleep as much as possible has been a long-standing and popular strategy to maximize the network lifetime (for example, the development of duty-cycle protocols [merlin2010duty, ye2006ultra, li2005energy]). Why optimize listening in contention resolution protocols if packets must receive messages (for other network applications)? Informally, there are two types of listening: listening to receive messages, and listening to execute a contention resolution protocol (which is what enables sending messages). To receive messages, packets do not need to listen in every time slot; there exist methods for optimizing this first type of listening. Although this is beyond the scope of this paper, as one example, in many wireless settings, a central base station can monitor the channel constantly and facilitate message exchange [feng2012survey].
We call a protocol fully energy-efficient if it is both sending efficient and listening efficient. By definition, such a protocol cannot have short feedback loops, since it can access the channel only rarely.
@@
Currently, the only known path to full energy efficiency is via explicit synchronization. This means that the model is extended so that packets can send synchronization messages to each other whenever they broadcast [BenderKoPe16, DeMarcoSt17, marco:time]. In [BenderKoPe16], these synchronization messages have size $\Theta(\ln N)$ bits each, which means that an arbitrary polynomial amount of communication can be expressed in a slot. In [DeMarcoSt17, marco:time], the synchronization messages are smaller ($O(1)$ bits), but some packets are permitted to stick around as “Good Samaritans” in order to serve as long-term coordinators (that send many messages via many broadcasts). We highlight that the model used in [DeMarcoSt17, marco:time] is directly comparable to our model. The language used in those works frames the problem as devices attempting to successfully send packets on a shared channel. In that language, the devices are allowed to remain in the system issuing additional coordination messages after successfully sending their packet.
Using $\Theta(\ln N)$-bit synchronization messages, Bender, Kopelowitz, Pettie, and Young [BenderKoPe16] give an algorithm with $\Theta(1)$ throughput and expected $O(\ln(\ln^{*}N))$ channel accesses per packet. Using Good-Samaritan packets, De Marco and Stachowiak [DeMarcoSt17] and De Marco, Kowalski, and Stachowiak [marco:time] provide a constant-throughput algorithm that is sending efficient ($O(\ln N)$ transmissions per packet) and conjecture that their techniques can be extended to achieve fully energy efficiency ($O(\operatorname{polylog}N)$ channel accesses per packet), with high probability and even without collision detection.
@@
We show that it is indeed possible to achieve constant throughput with full energy efficiency while being robust to adversarial noise. Moreover, these results hold in the standard ternary-feedback model, without requiring the addition of any sort of explicit synchronization.
@@
Our algorithm belongs to a natural family of multiplicative-weight-update algorithms (e.g., [ChangJiPe19, awerbuch:jamming, richa:jamming2, richa:jamming3, ogierman:competitive, richa:jamming4]). When a packet hears silence, it multiplicatively increases both its listening and sending probabilities. Conversely, when a packet hears noise, it multiplicatively decreases these probabilities. There are no control messages, no Good Samaritan packets, and no leaders elected.
@@
What makes our algorithm/analysis interesting is that we are able to support a multiplicative-weight-update framework while having each packet ‘cover its eyes’ almost all of the time. This is in stark contrast to prior constant-throughput algorithms, which adjust sending probabilities in every slot. Because each packet listens to so few slots, the different packets that are in the system at the same time may end up with very different perspectives on the world from each other. In order to analyze the ‘herd behavior’ of the packets in this potentially chaotic setting, substantially new techniques end up being required. These techniques are also what allow us to handle the additional chaos that adversarial jamming adds to the system.
1.1 Model
@@
A finite or infinite stream of indistinguishable packets arrives over time; the number of arrivals is unknown to the algorithm. Each packet must be sent on the multiple-access channel. Time is divided into synchronized slots, each of which is sufficiently large to send a single packet. An adversary (specified below) controls how many packets are injected into the system in each slot. When a packet succeeds, it departs the system. There is no universal numbering scheme for the slots; that is, there is no global clock from which a packet could infer the system lifetime or slot parity. Additionally, the packets do not receive any additional information about how many packets have arrived or will arrive. Instead, packets only receive information through the ternary feedback model.
@@
We now describe the ternary feedback model. In each time slot, each packet in the system can take one of three actions: (i) sleep, (ii) send, or (iii) listen to the channel. Packets that take actions (ii) or (iii) are said to access the channel. If no packets choose to send during a slot, then that (non-jammed) slot is empty/silent. A packet that listens during a slot (action iii) learns whether the slot was (0) empty, (1) successful, or (2+) noisy. A packet that sleeps during a slot (action i) learns nothing about the state of the slot. A packet that sends (action ii), either succeeds and leaves the system, or collides and remains in the system. [^2] For ease of presentation, in our algorithms, we say that a packet can listen and send simultaneously, but any packet that is sending actually does not need to listen to determine the state of the channel. If the packet is still in the system after sending in slot $t$, then slot $t$ was noisy.
@@
A basic metric: (overall) throughput.
The main objective of contention resolution is to optimize throughput, defined next. A slot is active if at least one packet is in the system during that slot; inactive slots can be ignored in our analysis. Similarly, a packet is active if has been injected into the system and has not yet departed. Without loss of generality, assume throughout that the first slot is active. Without jamming, the throughput at time $t$ is defined as $T_{t}/S_{t}$, where $T_{t}$ is the number of successful transmissions during slots $1,2,\ldots,t$, and $S_{t}$ is the number of active slots in the same time interval. [^3] By assumption that the first slot is active, we have $S_{t}\geq 1$ and hence throughput is always well defined. On a finite input, the (overall) throughput is defined with respect to the final active slot $t$; at this point, the overall throughput is $N/S$, where $N=T_{t}$ is the total number of packets and $S=S_{t}$ is the total number of active slots. Overall throughput is not well-defined on an infinite execution.
A deficiency of the throughput metric (when defined naïvely) is that even if an algorithm guarantees $\Theta(1)$ overall throughput, it is not possible to achieve $\Theta(1)$ throughput uniformly across time. For example, if there is a burst of $N$ packets at time 0, and $N$ is unknown to the algorithm, then the throughput will be 0 for a superconstant number of slots (e.g., see [willard:loglog, ChangKoPe19]), and this is provably unavoidable regardless of the backoff strategy being used. Thus (overall) throughput is only meaningful at the end of the execution, or at points in time where there are no packets in the system.
A stronger metric: implicit throughput [BenderKoKu20].
In order to support a meaningful notion of throughput, even at intermediate points in time, Bender, Kopelowitz, Kuszmaul, and Pettie [BenderKoKu20] propose a refined definition that they term “implicit throughput” [BenderKoKu20]. The implicit throughput at time $t$ is defined as $N_{t}/S_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$ and $S_{t}$ is the total number of active slots so far. [^4] Notice that while $N_{t}$ depends on the adversary, the number of active slots depends on the algorithm, whose goal it is to make slots inactive by completing packets.
@@
Since we ignore slots where no packets are in the system, the implicit throughput of our algorithm (see Section 3) is guaranteed (w.h.p.) to be $\Omega(1)$. Moreover, for every slot $t$ where the implicit throughput up to that slot is sufficiently small in $\Theta(1)$, the standard notion of throughput is also $\Theta(1)$.
Throughout the rest of the paper, we focus exclusively on implicit throughput. We should emphasize, however, that this only makes our results stronger—the results also imply the standard constant-throughput guarantees that one would normally strive for.
Extending to adversarial jamming.
We next extend the definitions of throughput and implicit throughput for the case of jamming following [BenderFiGi19]. An algorithm wastes a slot if that slot has silence or a collision, and throughput measures the fraction of slots that the algorithm could have used but instead wasted. Let $\mathcal{J}_{t}$ denote the number of jammed slots through slot $t$. Then, the throughput of an execution ending at time $t$ is defined to be $(T_{t}+\mathcal{J}_{t})/S_{t}$, and the implicit throughput at slot $t$ is defined to be $(N_{t}+\mathcal{J}_{t})/S_{t}$.
@@
[…]
Let $\delta$ be any lower bound on the implicit throughput of an algorithm; that is, suppose the algorithm achieves implicit throughput of at least $\delta$ at all times. Let $N_{t}$ and $S_{t}$ denote the total number of packet arrivals and active slots, respectively, at or before time $t$. Then $S_{t}\leq N_{t}/\delta$. Consequently: